Identify the following surfaces as an elliptical paraboloid, hyperbolic paraboloid, a hyperboloid of one sheet, a hyperboloid of two sheets, a cone, a circular cylinder, an elliptical cylinder, or a parabolic cylinder, and identify the axis of symmetry as the x-axis, the y-axis, or the z-axis.

a)−4x^2+9y^2−z^2=0
b)9y^2−5x^2−8z^2+1=0
c)5x^2+8y^2−z=0
d)6y^2+6z^2=1
e)−8x^2−4y^2+z^2=1
1
Expert's answer
2020-09-23T14:16:49-0400

a) "\u22124x^2+9y^2\u2212z^2 = 0" A cone with y-axis as the axis of symmetry

as y can be expressed in terms of x and z, 9y2 = 4x2 + z2



b) "9y^2\u22125x^2\u22128z^2+1=0" Hyperboloid of One Sheet with the y-axis as the axis of symmetry



c) "5x^2+8y^2\u2212z=0" Elliptic paraboloid with z-axis as the axis of symmetry

(z = 5x2 + 8y2)



d) "6y^2+6z^2=1" Elliptic cylinder along the plane yz parallel to x- axis i.e. the x-axis as the axis of symmetry



e) "\u22128x^2\u22124y^2+z^2=1" Hyperboloid of Two sheets with z-axis as the axis of symmetry